OpenStudy (anonymous):

Will give you a medal! Which is a prime number? A. 39 B. 49 C. 81 D. 89

OpenStudy (anonymous):

D

OpenStudy (anonymous):

89

OpenStudy (jadeishere):

89 A prime number is one that cannot be divided evenly by any number

OpenStudy (anonymous):

89 because it has factors of only 1 and itself

OpenStudy (anonymous):

89 is the prime number out of these

OpenStudy (jadeishere):

13 * 3 = 39 7 * 7 = 49 9 * 9 = 81 27 * 3 = 81

OpenStudy (anonymous):

Thankyou

OpenStudy (anonymous):

#toeasy!! xD

OpenStudy (anonymous):

hmm. yeah D

OpenStudy (texaschic101):

wow...you definitely got your answer

OpenStudy (anonymous):

A: 3x13 B: 7x7 C: 9x9 but D: the only possible one is 1x89

OpenStudy (anonymous):

Because the factors of 39 are 1,3,13,39 and the factors of49 are 1,7,and49 and the factors of 81 are 1,3,8,27, and 81 but 89 only have factors of 1 and 89

OpenStudy (anonymous):

Haha Thankyou guys! :)